# Splat-Nav: Safe Real-Time Robot Navigation in Gaussian Splatting Maps

### v1

@@ III-C Collision Detection
This proposition was stated by [15], but there was no proof. We write the proof explicitly here. The goal is to diagonalize both $\Sigma_{a}$ and $\Sigma_{b}$ in such a way that they share the same eigenbasis. In this way, the matrix inversion amounts to a reciprocal of the eigenvalues. We can find such an eigenbasis through the generalized eigenvalue problem, which solves the following system of equations
$\Sigma_{a}\phi_{i}=\lambda_{i}\Sigma_{b}\phi_{i}.$
These generalized eigenvalues and eigenvectors satisfy the identity
$\Sigma_{a}\phi=\Sigma_{b}\phi\Lambda.$
@@
To solve the generalized eigenvalue problem, we utilize the fact that $\Sigma_{a}$ and $\Sigma_{b}$ are symmetric, positive-definite, so the Cholesky decomposition of $\Sigma_{b}=LL^{T}$, where $L$ is lower triangular. This also means that we can represent $\Sigma_{a}$ by construction as
$\Sigma_{a}=LCL^{T}=LV\Lambda V^{T}L^{T},$
where $C$ is also symmetric, positive-definite, so its eigen-decomposition is $C=V\Lambda V^{T}$, where the eigenvectors $V$ are orthonormal $V^{T}V=I$. To retrieve the original eigenbasis $\phi$, we solve the triangular system $L^{T}\phi=V$. A consequence of this solution is that
$V^{T}V=\phi^{T}LL^{T}\phi=\phi^{T}\Sigma_{b}\phi=I.$
Moreover, this means that $\Sigma_{b}=\phi^{-T}\phi^{-1}$. We can show that $\Lambda$ and $\phi$ indeed solve the generalized eigenvalue problem through the following substitution
@@
As a result, we can write the matrix inversion in Theorem 1 as the following
@@
$\begin{split}\left[\frac{1}{1-s}\Sigma_{a}+\frac{1}{s}\Sigma_{b}\right]^{-1}&=\left[\frac{1}{1-s}\Sigma_{b}\phi\Lambda\phi^{T}\Sigma_{b}+\frac{1}{s}\Sigma_{b}\right]^{-1}\\ &=\left[\phi^{-T}\left(\frac{1}{1-s}\Lambda+\frac{1}{s}I\right)\phi^{-1}\right]^{-1}\\ &=\phi\left[\frac{1}{1-s}\Lambda+\frac{1}{s}I\right]^{-1}\phi^{T}\\ &=\phi\;\mathrm{\mathbf{diag}}\left(\frac{s(1-s)}{1+s(\lambda_{i}-1)}\right)\;\phi^{T}.\end{split}$
Multiplying the left and right by $\mu_{b}-\mu_{a}$ completes the proof. ∎
Corollary 1.
A finite number of samples for $s\in(0,1)$ will always find the presence of ellipsoid intersections. Hence, sampling instead of analytically solving Theorem 1 will still always yield safe behavior.
Proof.
@@
Using a finite number of samples, the test can fail (no sample points yield $K(s)>1$) or succeed (there is a sample point such that $K(s)>1$). Should the test succeed, by definition, there is no intersection. If it fails, there either exists an un-sampled point with $K(s)>1$ or such a point does not exist. If no such point exists, then the test remains correct in predicting collision. If such a point exists, it simply means our test predicted collision while in reality the ellipsoids are intersection-free. In other words, using finite samples leads to a safe, but conservative test. Due to the concavity of $K(s)$, more samples will asymptotically decrease the conservative-ness. ∎
The intersection test from Proposition 1 can now be parallelized by sampling many points between 0 and 1 and finding if any yield $K(s)>1$. This operation can also be parallelized across batches of ellipsoids, leading to massive computation speedups.
Corollary 2.
Scaling the ellipsoids $\mathcal{E}_{a},\mathcal{E}_{b}$ in Proposition 1 by a uniform positive scale factor $\alpha$ effectively scales $K(s)$ by a factor of $1/\alpha$.
In the case of uncertainty, either in the state of the agent or the confidence in the Gaussian Splatting representation, a margin of safety is needed in order for experiments outside of simulation to be correct. We can scale the representation and/or the robot body according to Corollary 2 to achieve the desired safety margins.
Corollary 3.
Gaussian Splatting parametrizes the covariance of the $j$th ellipsoid as $\Sigma_{j}=RSS^{T}R^{T}$, where $R^{T}$ is an orthonormal matrix that rotates the world coordinates into a body aligned frame, and $S$ is a diagonal matrix representing the axes lengths. Let $SS^{T}=\mathrm{\mathbf{diag}}\left(\lambda_{i}\right)$ and let $w=R^{T}(\mu_{\mathcal{R}}-\mu_{j})$. In addition, if we choose to parameterize our robot body as a sphere with covariance $\Sigma_{\mathcal{R}}=\kappa\mathcal{I}$ then the intersection test can be simplified to
Note that this test is faster to query than in Proposition 1 as there is no need to perform an eigenvalue decomposition. This again has computation speedups. For non-spherical robots, one can approximate it as a sphere by setting $\kappa$ to be the maximum eigenvalue of $\Sigma_{\mathcal{R}}$ (i.e., $\kappa=\lambda_{\rm max}(\Sigma_{\mathcal{R}})$).
III-D Finding a Small Test Set
Although performing the collision test in Corollary 3 is fast, it is difficult to naïvely scale to test all ellipsoids in the scene (which is often in the 10’s or 100’s of thousands for GSplat maps) while maintaining real-time performance in a planner. As a result, we rely on the cost-effectiveness of KD-tree queries in our proposed algorithm to quickly converge on a small set of ellipsoids to test.
Input: Query point $x$, robot’s largest eigenvalue $\kappa$, ellipsoids $\mathcal{G}=\{\mu_{i},\Sigma_{i}\}_{i=1}^{N}$;
Output: Minimal Collision Set $\mathcal{G}_{x}\subseteq\mathcal{G}$;
$\lambda_{\rm max}(\mathcal{I})\triangleq\max_{i\in\mathcal{I}}\lambda_{\rm max}(\Sigma_{i})$;
KD $\leftarrow$ KDTree($\{\mu_{i}\}$);
$\mathcal{I}_{0}\leftarrow\{1,\ldots,N\}$;
$\mathcal{I}_{1}\leftarrow$ KD.QueryBallPoint($x$, $\lambda_{\max}(\mathcal{I}_{0})+\kappa$);
while $\lambda_{\max}(\mathcal{G}_{1})<\lambda_{\max}(\mathcal{I}_{0})$ or $\mathcal{I}_{1}\neq\varnothing$ do
 $\mathcal{I}_{0}\leftarrow\mathcal{I}_{1}$ ;
 $\mathcal{I}_{1}\leftarrow$ KD.QueryBallPoint($x$, $\lambda_{\max}(\mathcal{I}_{0})+\kappa$);
end while
$\mathcal{G}_{x}=\{\mu_{i},\Sigma_{i}\}_{i\in\mathcal{I}_{1}}$;
Algorithm 1 Minimal Collision Test Set
The minimal collision test set $G_{x}$ is a ball centered at a test point $x$ with a radius that guarantees the ball will contain the center of the largest ellipsoid (with largest eigenvalue $\lambda_{\max}$) whose circumscribed circle intersects with the robot body centered at $x$. We know for certain that ellipsoids outside of this ball cannot intersect with the robot body. We also know that necessarily, this ball must have radius at most $\lambda_{\rm max}+\kappa$, meaning the center of this ellipsoid lies on the surface of the ball. To find this minimal test set, we propose using Algorithm 1. We stop iterating when the largest eigenvalue between successive queries to the KD-tree does not change, meaning the ellipsoid corresponding with the largest eigenvalue is the largest ellipsoid that could intersect our robot. The algorithm will only find the minimal set if the center associated with $\lambda_{\rm max}$ lies on the surface of the ball. However, the gap between the converged set and the minimal set is relatively small.
III-E Supporting Hyperplanes
We like to again emphasize that having convex primitives (ellipsoids) as an environment representation facilitates development of interpretable algorithms. In this work, it is the ability to create polytopes that describe the safe regions of space through the use of supporting hyperplanes. Our contribution is inspired by [27]. However, there are two key differences. Liu et al. [27] assumes a point cloud as the environment, and requires an initial, safe path to create the polytopes. On the other hand, the safe polytopes generated using a GSplat are generalizable. Assuming a valid safe polytope exists, they can be generated by any sequence of points regardless of whether the points are in collision.
@@ III-E Supporting Hyperplanes
Given a test point $x^{*}$ (i.e., robot position $\mu_{\mathcal{R}}$) and a collision test set $\mathcal{G}^{*}$, for every ellipsoid in $\mathcal{G}^{*}$, a supporting hyperplane to the $j$-th ellipsoid (considering the extent of the robot body) derived from Proposition 1 is given by
$\underbrace{\Delta^{T}_{j}Q_{j}}_{a_{j}}x\geq\underbrace{\frac{k_{j}^{*}}{1+\epsilon}+\Delta^{T}_{j}Q_{j}\mu_{j}}_{b_{j}},$
for any $\epsilon>0$ where $\Delta_{j}=x^{*}-\mu_{j}$, $s^{*}=\operatorname*{arg\,max}_{s\in(0,1)}K(s)$, $Q_{j}=\phi\;\mathrm{\mathbf{diag}}\left(\frac{s^{*}(1-s^{*})}{1+s^{*}(\lambda_{i}-1)}\right)\phi^{T}$, and $(k_{j}^{*})^{2}=K(s^{*})=\Delta^{T}_{j}Q_{j}\Delta_{j}>0$. Note that $Q_{j}$ is always symmetric, positive-definite. By stacking the hyperplane constraints ($a_{j},b_{j}$), we arrive at a polytope $Ax\geq b$ that is guaranteed to be safe.
Proof.
@@
From Proposition 1, we know that we must find an $x$ such that $K(s)>1$ for all ellipsoids in the test set $\mathcal{G}^{*}$. We will approximate this safe set as a polytope. Given a test point $x^{*}$ and the $j$th ellipsoid in the collision test set $\mathcal{G}^{*}$, we can use our collision test (Corollary 3) to derive these polytopes. Notice in (3) that $\mu_{a}$ represents the position of the robot (i.e., the test point $x^{*}$) and $\mu_{b}$ is the centroid of ellipsoid $j$. Once we find the value $s^{*}$ that maximizes $K(s)$, the safety test is a quadratic constraint of the test point $x^{*}$, namely $\Delta_{j}^{T}Q_{j}\Delta_{j}>1$.
@@
To derive the supporting hyperplane for the safety check, we will first find the point on the ellipsoid and then linearize our test about this point. Let $f_{j}(x)=(x-\mu_{j})^{T}Q_{j}(x-\mu_{j})$. Note that $f_{j}(x^{*})=\Delta^{T}_{j}Q_{j}\Delta_{j}=(k_{j}^{*})^{2}$. We can immediately see that point $x_{0}=\mu_{j}+\frac{1+\epsilon}{k_{j}^{*}}\Delta_{j}$ will be outside of the ellipsoid for any $\epsilon>0$ since $f_{j}(x_{0})=(x_{0}-\mu_{j})^{T}_{j}Q_{j}(x_{0}-\mu_{j})=\frac{(1+\epsilon)^{2}}{(k_{j}^{*})^{2}}\Delta_{j}^{T}Q_{j}\Delta_{j}=(1+\epsilon)^{2}>1$. Note that this point $x_{0}$ lies on the ray starting from the center of the ellipsoid $\mu_{j}$ and passing through the center of the robot at $x^{*}$. We then linearize the constraint $f_{j}(x)$ about $x_{0}$. Taking the derivative yields $\frac{df_{j}}{dx}(x)=2(x-\mu_{j})^{T}Q_{j}$. The linear approximation of the constraint is then $f_{j}(x)\approx f_{j}(x_{0})+\left[\frac{df_{j}}{dx}(x_{0})\right](x-x_{0})=(1+\epsilon)^{2}+2(x_{0}-\mu_{j})^{T}Q_{j}(x-x_{0})\geq 1$. Plugging in $x_{0}=\mu_{j}+\frac{1+\epsilon}{k_{j}^{*}}\Delta_{j}$ and simplifying yields the given expression $\Delta^{T}_{j}Q_{j}x\geq\frac{k_{j}^{*}}{1+\epsilon}+\Delta^{T}_{j}Q_{j}\mu_{j}$.

### v2

@@ Appendix A Proof of Proposition 1
This proposition was stated by [59] without proof. We provide the proof here. The goal is to diagonalize both $\Sigma_{a}$ and $\Sigma_{b}$ in such a way that they share the same eigenbasis. In this way, the matrix inversion amounts to a reciprocal of the eigenvalues. We can find such an eigenbasis through the generalized eigenvalue problem, which solves the following system of equations $\Sigma_{a}\phi_{i}=\lambda_{i}\Sigma_{b}\phi_{i}.$} These generalized eigenvalues and eigenvectors satisfy the identity $\Sigma_{a}\phi=\Sigma_{b}\phi\Lambda.$}
@@
To solve the generalized eigenvalue problem, we utilize the fact that $\Sigma_{a}$ and $\Sigma_{b}$ are symmetric, positive-definite, so the Cholesky decomposition of $\Sigma_{b}=LL^{T}$, where $L$ is lower triangular. This also means that we can represent $\Sigma_{a}$ by construction as: $\Sigma_{a}=LCL^{T}=LV\Lambda V^{T}L^{T},$} where $C$ is also symmetric, positive-definite, so its eigen-decomposition is $C=V\Lambda V^{T}$, where the eigenvectors $V$ are orthonormal $V^{T}V=I$. To retrieve the original eigenbasis $\phi$, we solve the triangular system $L^{T}\phi=V$. A consequence of this solution is that: $V^{T}V=\phi^{T}LL^{T}\phi=\phi^{T}\Sigma_{b}\phi=I.$} Moreover, this means that $\Sigma_{b}=\phi^{-T}\phi^{-1}$. We can show that $\Lambda$ and $\phi$ indeed solve the generalized eigenvalue problem through the following substitution:
@@
As a result, we can write the matrix inversion in Theorem 1 as the following:
@@
$\begin{split}\left[\frac{1}{1-s}\Sigma_{a}+\frac{1}{s}\Sigma_{b}\right]^{-1}&=\left[\frac{1}{1-s}\Sigma_{b}\phi\Lambda\phi^{T}\Sigma_{b}+\frac{1}{s}\Sigma_{b}\right]^{-1}\\ &=\left[\phi^{-T}\left(\frac{1}{1-s}\Lambda+\frac{1}{s}I\right)\phi^{-1}\right]^{-1}\\ &=\phi\;\mathrm{\mathbf{diag}}\left(\frac{s(1-s)}{1+s(\lambda_{i}-1)}\right)\;\phi^{T}.\end{split}$
Multiplying the left and right by $\mu_{b}-\mu_{a}$ completes the proof.
Appendix B Proof of Corollary 1
@@
Using a finite number of samples, the test can fail (no sample points yield $K(s)>1$) or succeed (there is a sample point such that $K(s)>1$). Should the test succeed, by definition, there is no intersection. If it fails, there either exists an un-sampled point with $K(s)>1$ or such a point does not exist. If no such point exists, then the test remains correct in predicting collision. If such a point exists, it simply means our test predicted collision while in reality the ellipsoids are intersection-free. In other words, using finite samples leads to a safe, but conservative test. Due to the concavity of $K(s)$, more samples will asymptotically decrease the conservativeness. In fact, due to the monotonicity of the derivative of $K(s)$ and the bounds on $s$, even performing iterations of bisection search will yield exponential convergence.
Appendix C Ellipsoid-to-Ellipsoid Intersection Tests
In the subsequent discussion, we examine existing algorithms for detecting if a pair of ellipsoids intersect. Recall that we have the original test in Proposition 1 that involves solving the generalized eigenvalue problem and the sampling-based approximation in Corollary 1, with guarantees on safety due to function convexity.
A Test Based on Sherman-Morrison formula
We introduce an alternative approach for solving Theorem 1 that neither requires solving a generalized eigenvalue problem nor explicitly computing the inverse of a matrix. Given the representation of the ellipsoids: $\Sigma_{a}=R_{a}S_{a}S_{a}^{T}R_{a}^{T}$} and $\Sigma_{b}=R_{b}S_{b}S_{b}^{T}R_{b}^{T}$}, let $S_{a}=\mathrm{\mathbf{diag}}(\sqrt{\lambda_{a,(1)}},\ldots,\sqrt{\lambda_{a,(3)}})$}, with $S_{b}$} defined similarly.
@@ A Test Based on Sherman-Morrison formula
The concave function $K(s)$ in the intersection test given in Theorem 1 can be simplified to:
$K(s)\coloneqq(\mu_{a}-\mu_{b})^{T}V_{3}(s)(\mu_{a}-\mu_{b}),$ (9)
where $V_{3}:\mathbb{R}\rightarrow\mathbb{R}^{3\times 3}$} represents a matrix-valued function of $s$. For simplicity, we present $V_{3}(s)$} as a composition of the following functions of $s$: $\forall i\in\{1,2,3\}$},
$V_{i}(s)$ $\coloneqq V_{i-1}(s)-\frac{V_{i-1}(s)d_{b,(i)}r_{b,(i)}r_{b,(i)}^{T}V_{i-1}(s)}{1+d_{b,(i)}r_{b,(i)}^{T}V_{i-1}(s)r_{b,(i)}},$
with $V_{0}(s)\coloneqq R_{a}\Lambda_{a}(s)R_{a}^{T},$} where $\Lambda_{a}:\mathbb{R}\rightarrow\mathbb{R}^{3\times 3}$} yields a diagonal matrix, whose $i$th diagonal element is given by $\frac{(1-s)}{\lambda_{a,(i)}}$, $r_{b,(i)}\in\mathbb{R}^{3}$} denotes the $i$th column of $R_{b}$, and $d_{b,(i)}\coloneqq\frac{\lambda_{b,(i)}}{s}$}.
We omit the proof as it follows from the iterative application of the Sherman-Morrison formula [60].
B Results
We compare the methods in [61] to the collision-detection algorithms presented in Proposition 1, Corollary 1, and Proposition 3. We did not compare against the method in [9], as well as optimization-based methods that require the computation of a separating hyperplane, since these methods do not scale efficiently to dense scenes with many Gaussians, which arise in Gaussian Splatting Maps. Figure 12 shows the results, where we see that Corollary 1 is the fastest across the full range of map densities.
Figure 12: Comparison of mean compute times and performance gap for various ellipsoid-ellipsoid collision detection algorithms, each executed on 100 different sets of Gaussians per fixed number of Gaussians.
Appendix D Minimal Collision Test Set
The minimal collision test set $G_{x}$ is a ball centered at a test point $x$ with a radius that guarantees the ball will contain the center of the largest ellipsoid (with largest eigenvalue $\lambda_{\max}$) whose circumscribed circle intersects with the robot body centered at $x$. We note that ellipsoids outside of this ball cannot intersect with the robot body. Further, this ball must have radius at most $\lambda_{\rm max}+\kappa$}, meaning the center of this ellipsoid lies on the surface of the ball. We stop iterating when the largest eigenvalue between successive queries to the K-D tree does not change, meaning the ellipsoid corresponding with the largest eigenvalue is the largest ellipsoid that could intersect our robot. The algorithm will only find the minimal set if the center associated with $\lambda_{\rm max}$ lies on the surface of the ball. However, the gap between the converged set and the minimal set is relatively small.
Appendix E Proof of Proposition 2
@@
From Proposition 2, we know that we must find an $x$ such that $K(s)>1$ for all ellipsoids in the test set $\mathcal{G}^{*}$. We will approximate this safe set as a polytope. Given a test point $x^{*}$ and the $j$th ellipsoid in the collision test set $\mathcal{G}^{*}$, we can use our collision test (Corollary 2) to derive these polytopes. Notice that $\mu_{a}$ represents the position of the robot (i.e., the test point $x^{*}$) and $\mu_{b}$ is the centroid of ellipsoid $j$. For any value of $\bar{s}\in(0,1)$}, the safety test is a quadratic constraint of the test point $x^{*}$, namely $\Delta_{j}^{T}Q_{j}\Delta_{j}>1$.
@@
To derive the supporting hyperplane for the safety check, we will first find the point on the ellipsoid and then linearize our test about this point. Let $f_{j}(x)=(x-\mu_{j})^{T}Q_{j}(x-\mu_{j})$. Note that $f_{j}(x^{*})=\Delta^{T}_{j}Q_{j}\Delta_{j}=k_{j}^{2}$. We can immediately see that point $x_{0}=\mu_{j}+\frac{1+\epsilon}{k_{j}}\Delta_{j}$ will be outside of the ellipsoid for any $\epsilon>0$ since:
$f_{j}(x_{0})$ $=\frac{(1+\epsilon)^{2}}{k_{j}^{2}}\Delta_{j}^{T}Q_{j}\Delta_{j}=(1+\epsilon)^{2}>1.$ (10)
Note that this point $x_{0}$ lies on the ray starting from the center of the ellipsoid $\mu_{j}$ and passing through the center of the robot at $x^{*}$. We then linearize the constraint $f_{j}(x)$ about $x_{0}$. Taking the derivative yields $\frac{df_{j}}{dx}(x)=2(x-\mu_{j})^{T}Q_{j}$. The linear approximation of the constraint is then:
$f_{j}(x)$ $\approx f_{j}(x_{0})+\left[\frac{df_{j}}{dx}(x_{0})\right](x-x_{0})\geq 1.$ (11)
Plugging in $x_{0}=\mu_{j}+\frac{1+\epsilon}{k_{j}}\Delta_{j}$ and simplifying yields the given expression $\Delta^{T}_{j}Q_{j}x\geq{(1+\epsilon)}{k_{j}}+\Delta^{T}_{j}Q_{j}\mu_{j}$.